# Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

### v1

@@ IV-A1 Exact evaluation
Exact evaluation is based on a simple idea: counting and heaps [37]. Assume that the sample data of items has been already arranged from largest to smallest. Solution $S=[s_{1},s_{2},\ldots,s_{m}]$ is under evaluated. The details of exact evaluation are listed as follows:
@@
    Then push neighbors of the popped $D$ whose sum are smaller than that of $D$ into $Q$, listed as follows:
@@
We have the following theorem,
@@
Detailed proofs can be found in Appendix 2. If the ratio of the number to the total number of combinations is greater than $1-P_{0}$, then we determine that the solution is infeasible.
@@ IV-A2 Accelerated Monte Carlo evaluation
Monte Carlo simulation is a method that uses random numbers to solve many computational problems [38]. In this work, we focus on situations where the knapsack constraints $W$ are tight at a high confidence level $P_{0}$, which means feasible solutions account for a very small proportion of the entire solution space of the problem. This observation motivates us that the evaluation procedure can be greatly accelerated if most of the infeasible solutions can be quickly eliminated. Specifically, assuming that the sample data of the items has been sorted from largest to smallest, we can check the $l^{th}$ largest data combination $d_{1l},d_{2l},...,d_{ml}$ for the combination of items to be solved, i.e.,$D_{l}=\sum_{i=1}^{m}d_{il}$, by comparing with the weight threshold $W$ to directly filter out most of the infeasible solutions. Therefore, the evaluations times required during the algorithm operation can be significantly reduced. The process can be summarized as follows.
  • Construct a filtered list according to the minimum prime factorization for $K=L^{m}\times(1-P_{0})$.
  • Only a limited size of combinations will be selected to construct the filtered list $R\ll m!$;
  • Calculate the sum of the data of each item of the screening list $R$, and check whether there is at least one combination satisfying $\sum_{i=1}^{m}d_{il}\geq W$. If not, the solution will be evaluated by Monte Carlo simulation.
@@ IV-B Surrogate Weight Design
It is not feasible to directly obtain the cost-effectiveness of each item in a data-driven scenario, which impedes the algorithm’s ability to search for better solutions. To overcome this challenge, we construct a surrogate cost-effectiveness based on the statistic of sample data. Specifically, we calculate the sample mean and standard deviation of the item weights, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$. Then item weights are represented by the weighted sum of statistical characteristics,
@@
where $\lambda$ is the weighted parameter. This approach provides a statistical measure for guiding the search process towards optimal solutions.
@@ IV-C Local Search Framework
Based on the previously data-driven solution evaluation method and surrogate weight design, we now introduce the overall algorithmic framework of DDALS. Local search is a widely used heuristic search framework for efficiently solving various combinatorial optimization problems [12, 39]. Thus, our algorithm adopts the framework of local search algorithm, which can adaptively jump between small and large regions, achieving efficient search for solutions. The block diagram of DDALS shown in Figure 3 . DDALS is composed of the following five components: constructive procedure (CP) for solution initialization, a local swap search (LSS) operator for improving the current solution, a degradation (Degrade) operator for escaping local optima, and a further swap search (FSS) operator for searching in a larger feasible region. Since limited sample data cannot exactly reflect the true distribution, we additionally incorporate a solution feasibility enhancement (SFE) to improve the feasibility of the output.
@@
Input: A DDCCMCKP benchmark; Maximum iteration times, $MaxIter$; weighting factor, $\lambda$
@@ 11 Return $S_{opt}$;
Algorithm 1 illustrates the framework of DDALS. The inputs of DDALS include the benchmark, maximum iteration times $MaxIter$, and a user-defined weighting factor $\lambda$. We first initialize the solution $S_{cp}$ using CP (line 1), then iteratively improve the quality of solutions. At iteration $t$, LSS attempts to obtain a better feasible solution by operating on $S_{t}$ (line 3). If LSS finds an improvement, the optimal solution $S^{*}$ is updated (line 4-5). The current solution $S_{t}$ will be degraded by randomly selecting an item for replacement to escape a local optimum (line 6). After iterations end, FSS is employed on $S_{opt}$ to search for a better solution (line 8). The final solution $S_{opt}$ will be output by SFE.
The detailed designs of the components are list as follows:
@@ IV-C1 Constructive procedure
The initialization procedure is based on greedy strategy [7]. Utilizing the item weights defined in Definition 3, we calculate the cost-effective variables $u_{ij}=v_{ij}/\tilde{w}_{ij}$ for each item $j\in N_{i}$, then sort the items in descending order of $u_{ij}$ for each class $N_{i}$. $S_{cp}$ is constructed by selecting the item with the largest $u_{ij}$ in each class. If the current $S_{t}$ is infeasible, the item which has the largest $\tilde{w}_{i_{0}j}$ in $S_{t}$ is replaced by an item with a smaller weight in the same class $i_{0}$. The procedure iterates until a feasible $S_{cp}$ is obtained. If all items in class $i_{0}$ are tried but still infeasible,$S_{cp}$ will be constructed by the smallest $\tilde{w}_{ij}$ in each class.
@@ IV-C2 Local swap search
LSS applies a swapping strategy for picked items to improve the current solution $S_{t}$. If the obtained solution is feasible, each swap is authorized. LSS will try all the items with lower cost in each class $N_{i}$ and selects the item with the lowest cost from all authorized swaps.
@@ IV-C3 Degrade
Degrade randomly replaces an item $j_{i_{0}}$ of class $i_{0}$ in the current solution $S_{t}$. If an infeasible solution is generated, it randomly selects a lighter item in $i_{0}$. If there is no suitable item in the $i_{0}$, it randomly switches to another class and tries again until a new feasible solution $S_{t}$ is obtained.
@@ IV-C4 Further swap search
FSS iterates through all class combinations $(i_{0},i_{1})$ of the current optimal solution $S_{t}$ to find a better and feasible solution. If it finds a pair of items with a smaller cost, $S_{t}$ will be updated. The algorithm ends when all pairs of items have been searched.
IV-C5 Solution feasibility enhancement
In practice, due to the demand for high confidence level $P_{0}$, we tend to minimize the cost of the solution while ensuring its feasibility, rather than blindly pursuing the minimization of the cost, leading to solutions that do not satisfy constraints, which renders them useless in practice. It is a crucial requirement in many practical applications. However, it is difficult to guarantee the real feasibility of the optimal solution due to limited sample size. Note that the sample size to achieve smaller errors for arbitrarily distributed random variables will be very large (See Appendix 3 for theoretical analysis between estimation error and sample size). Therefore, we design three filtering variants to enhance the feasibility of output solution, while not increasing the cost too much. In the iterations of DDALS algorithm, two list of feasible solutions are recorded, each with a capacity of 30. The first records 30 feasible solutions with the lowest cost encountered so far, named Cost List. The second records 30 solutions with the largest confidence level encountered so far, named MC List.
  • Original: Output the optimal solution of FSS.
  • Variant 1: The minimum-cost solution with a higher confidence level than $P_{0}$. For example, it should satisfy $P\geq 0.995$ if $P_{0}=0.99$.
  • Variant 2: All solutions will be scored according to the order in the list they are in. Each candidate solution has two scores, which will be added together. Choose the solution with the highest score as the final output.
  • Variant 3: Take the first 10 solution (if the amount is larger than 10) in Cost List as the final output and choose one manually.
@@ IV-D1 Exact evaluation
First, if the solution can be directly determined infeasible in the fast screening phase, the time complexity takes $O(1)$. Otherwise, in order to verify whether it is feasible, the maximum number of iterations that the algorithm needs to perform is $T=L^{m}(1-P_{0})-l_{min}^{m}$, where $l_{min}$ is the index that the combination of $l_{min}$ samples of solution is larger than $T_{\max}$. The maximum heap size is $k_{0}+(m-1)T=\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+(m-1)L^{m}(1-P_{0})-(m-1)l_{min}^{m}=O(mT)$. Since at each iteration it needs to sort the heap once, the total number of times is $O(\log(mT))=O(\log m+\log T)=O(\log T),m\ll L$. Thus, the total time complexity of the filtered algorithm is $O(mL\log L+\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+T\log T)=O(\max\{T\log T,ml_{min}^{m-1},m^{2}l_{min}^{m-2},...,ml_{min}\})$.
@@ IV-D2 Accelerated Monte Carlo evaluation
The complexity of the accelerated Monte Carlo evaluation is $O((|G|+\alpha\times MonteCarlo)\times Eval)$, where $MonteCarlo$ denotes the set number of Monte Carlo simulations and $|G|$ denotes the size of fast screening combination. If the solution is an infeasible solution determined by fast screening, $\alpha=0$, otherwise $\alpha=1$.
@@ IV-D3 DDALS
  • CP includes sorting and solution evaluation. Up to $m\times N$ times of evaluation are required. The complexity is $O(Eval\times mN)$, where $N=\max|N_{i}|$.
  • Similar to CP, LSS require up to $m\times N$ times of evaluation to search a possible better solution. The complexity is $O(Eval\times mN)$.
  • The worst case is that only one combination of items is feasible in the whole solution space. The complexity of Degrade is $O(Eval\times mN)$.
  • FSS iterates through a maximum of combinations of classes $\binom{m}{2}\times N^{2}$. The complexity of FSS is $O(Eval\times m(m-1)N^{2}/2)=O(Eval\times m^{2}N^{2})$.
Note that SFH only takes $O(1)$. Therefore, the total time complexity of DDALS is $O(Eval\times mN+MaxIter\times(Eval\times mN+Eval\times mN)+Eval\times N^{2}m^{2}+1)=O(Eval\times\max\{N^{2}m^{2},MaxIter\times mN\})$.
@@ V Benchmark Design
Since DDCCMCKP is a new problem that has not been solved before, it is necessary to design benchmarks to evaluate the performance of algorithms in different scenarios. Therefore, we designed two sets of benchmarks to cover a wide range of problem sizes and complexities, each of which consists of multiple benchmarks of different scales, in order to comprehensively evaluate the algorithm’s performance in different scenarios.
@@
The first set of benchmarks aims to simulate the case where the delay of items follows different common continuous distributions, named LAB. The distribution of each item is randomly chosen from the set of uniform, truncated normal, fatigue-life, bimodal and gamma distributions. The mean of items is uniformly distributed in $U(2,8)$, i.e., $\mu\sim U(2,8)$, and the variance is uniformly distributed in $U(1,19)$, i.e., $\sigma^{2}\sim U(1,19)$, where $U(2,8)$ denotes a uniform distribution over the interval [2,8]. In the case of the gamma distribution, the mean value is $\mu\sim U(0.5,2.5)$ and variance $\sigma^{2}\sim U(0.05,0.625)$. The cost, representing the value of taking the item, is generated randomly between $U(1,10)$, corresponding to uncorrelated situation in classic MCKP benchmark[26], which is a widely used benchmark in the literature. The curve of probability density function of different distributions of the items involved in LAB is shown in Figure.4a.
@@
The second set of benchmarks simulates the actual business requirements of a telecommunication company in the 5G E2E network, denoted as APP. In APP, the distribution of each item is also randomly chosen in the set of uniform, truncated normal, fatigue-life, bimodal and gamma distribution, which is similar to the LAB case. However, in the 5G E2E network, each data package transmission is restricted to a time window with a fixed size, indicating that packets must be transmitted successfully within a set interval. The difference between APP and LAB lies in the introduction of the retransmission mechanism [17]. This mechanism implies that each item in all classes has a probability of being retransmitted when a transmission failure occurs. In the APP, the first data package transmission attempt must be completed within $(0,10]$ ms with a success probability of 0.9. If the first attempt fails, a second attempt occurs with a delay of $(10,20]$ ms and a success probability of 0.9. Therefore, the delay of one transmission ranges from $(10,20]$ ms with a probability of $0.1\times 0.9=0.09$. The same applies to the third and fourth retransmissions. The transmission fails if an item requires transmitting data more than four times, which limits the maximum number of retransmissions. Consequently, the time delay for an item to transmit a data package belongs to the intervals $(0,10],(10,20],(20,30],(30,40]$ with probabilities of $0.9,0.09,0.009,0.001$, respectively. The distribution of items involved in the APP is presented in Figure 4b. The cost generation formula for items is $c_{ij}=10/(\mu_{ij}+\sigma_{ij})\times U(0.8,1.2)$. The purpose is to make the cost-effectiveness decrease as the delay decreases, thus simulating the marginal diminishing effect of cost-effectiveness in practical application.
For each benchmark in the LAB and APP, we design small-scale and large-scale problems with different parameter settings, as shown in Table I and II respectively. The scale of benchmarks, i.e., $m$ and $N$, is set similar to [26]. The sample sizes $L$ are set 30 for small-scale benchmarks and 500 for large-scale benchmarks. In Table I, ”{11,14}” of $T_{\max}^{LAB}$ means that $T_{\max}$ is set 11 or 14 as two different settings in the experiments of ss1.
TABLE I: Small-scale benchmarks of DDCCMCKP
Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$
@@
TABLE II: Large-scale benchmarks of DDCCMCKP
Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$

### v2

@@ IV-A1 Exact evaluation
Exact evaluation is based on counting and heaps [34]. Assume that the sample data of each item is already arranged from largest to smallest, and we have a solution $S=[s_{1},s_{2},\ldots,s_{m}]$ under evaluated. The details of the algorithm are listed as follows:
@@
    Then push neighbors of the popped $D$ whose sum is smaller than that of $D$ into $Q$, listed as follows:
@@
More importantly, we have the following theorem,
@@
Detailed proofs can be found online in the supplementary material. If the ratio of $k^{th}$ to the total number of combinations exceeds $1-P_{0}$, we conclude that the solution is infeasible.
@@ IV-A2 Accelerated Monte Carlo evaluation
Monte Carlo simulation uses random numbers to solve computational problems [35, 36]. In this work, our focus is on scenarios where the knapsack constraints $W$ are tightly defined at a high confidence level $P_{0}$, indicating that feasible solutions constitute a small proportion of the entire solution space. It motivates us that the evaluation procedure can be greatly accelerated if most of the infeasible solutions can be quickly recognized. Assuming that the sample data of items has been sorted in descending order, we check the $l^{th}$ largest data combination $d_{1l},d_{2l},...,d_{ml}$, compare them with the weight threshold $W$, thereby achieving the efficient screening of infeasible solutions. The process is summarized as follows.
  • Construct a screening list according to the minimum prime factorization for $K=L^{m}\times(1-P_{0})$.
  • A limited number of combinations are selected to construct the screening list $R\ll m!$;
  • Calculate the sum of the data of each item of the screening list $R$, and check whether there is at least one combination satisfying $\sum_{i=1}^{m}d_{il}\geq W$. If not, $S$ is evaluated by Monte Carlo simulation.
@@ IV-B Surrogate Weight Design
It is infeasible to directly obtain the cost-effectiveness of each item in a data-driven scenario. To overcome this challenge, we construct the surrogate weight of item $\tilde{w}_{ij}$ based on sample data. The sample mean and sample standard deviation of item are calculated, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$. The item weights are then calculated by the weighted sum of $\mu_{ij}$ and $\sigma^{2}_{ij}$,
@@
where $\lambda$ is the weighted parameter. This approach provides a statistical measure for constructing the surrogate weight from sample data.
@@ IV-C Local Search Framework
Local search is a widely used heuristic search framework for solving various combinatorial optimization problems [12, 37]. We integrate local search with data-driven evaluation approach and surrogate weight design, leading to DDALS. The block diagram is illustrated in Figure 3. DDALS comprises five components: constructive procedure (CP) for solution initialization, a local swap search (LSS) operator for improving the current solution, a degradation (Degrade) operator for escaping local optima, and a further swap search (FSS) operator for searching in a larger region. Considering that limited sample data can not precisely mirror the true distribution, we additionally incorporate a solution feasibility enhancement (SFE) to enhance the feasibility of the output solutions.
@@
Input: Problem intance; maximum iteration times, $MaxIter$; weighting factor, $\lambda$
@@ 11 Return $S_{opt}$;
Algorithm 1 illustrates the framework of DDALS. The inputs for DDALS include the problem instances, maximum iteration times $MaxIter$, and a pre-defined weighting factor $\lambda$. DDALS first initializes the solution $S_{cp}$ using CP (line 1) and iteratively enhances the quality of solutions. At iteration $t$, LSS attempts to improve the feasible solution by operating on $S_{t}$ (line 3). If LSS identifies an improvement, it updates the optimal solution $S^{*}$ (lines 4-5). The current solution $S_{t}$ is degraded by randomly selecting an item for replacement to escape a local optimum (line 6). Once iterations end, FSS is applied to search for a better solution (line 8). The final solution $S_{opt}$ is output by SFE. The detailed designs of these components are list as follows:
@@ IV-C1 Constructive procedure
The initialization procedure is based on greedy strategy [7]. Utilizing the surrogate weights defined in Definition 3, we calculate the cost-effectiveness variables $u_{ij}=v_{ij}/\tilde{w}_{ij}$ for each item $j\in N_{i}$, then sort the items in descending order of $u_{ij}$ for each class $N_{i}$. $S_{cp}$ is constructed by selecting the item with the largest $u_{ij}$ in each class. If the current $S_{t}$ is infeasible, the item which has the largest $\tilde{w}_{i_{0}j}$ in $S_{t}$ is replaced by the one with a smaller weight in the same class $i_{0}$. The procedure iterates until feasible $S_{cp}$ is obtained. If all items in class $i_{0}$ have been tried but $S_{cp}$ is still infeasible, $S_{cp}$ will be constructed by the items with the smallest surrogate weights.
@@ IV-C2 Local swap search
LSS applies a swapping strategy to improve the current solution $S_{t}$. If the obtained solution is feasible, the swap operation is authorized. LSS aims to select items with the smallest cost while ensuring that the constructed solution is feasible.
@@ IV-C3 Degrade
Degrade randomly replaces an item $j_{i_{0}}$ of class $i_{0}$ in the current solution $S_{t}$. If an infeasible solution is constructed, it randomly selects an item with smaller surrogate weight in class $i_{0}$. If there is no suitable item, Degrade randomly selects another class and tries again until a new feasible solution $S_{t}$ is obtained.
@@ IV-C4 Further swap search
FSS iterates through all the item combinations $(i_{0},i_{1})$ from two different classes of the current optimal solution $S_{t}$. When a pair of items with smaller total cost is found, $S_{t}$ is updated. The procedure terminates when all possible pairs of items have been tried.
IV-C5 Solution feasibility enhancement (SFE)
In practical scenarios, we aim to minimize costs while ensuring that the solution is feasible. Notably, achieving smaller errors for arbitrarily distributed random variables necessitates an exceedingly large sample size (see the online supplementary material for theoretical analysis). Due to limited sample size, instead of blindly minimizing costs, by allowing for a certain margin of error, the probability of obtaining a solution that is truly feasible in practical scenarios is larger. Therefore, we designed the SFE. When DDALS starts running, SFE maintains two lists, each recording up to 30 feasible solutions. The first list is called MC list, which records the 30 solutions with the highest confidence levels obtained during the DDALS search process. The second list is called Cost list, which records the 30 lowest-cost feasible solutions obtained throughout the whole operation of DDALS.
  • Original: DDALS outputs the optimal solution without enhancement.
  • Variant 1: The minimum-cost solution with a higher confidence level. For example, $S_{t}$ should satisfy that the confidence level is greater than 0.995 when $P_{0}=0.99$.
  • Variant 2: All solutions are scored within their respective lists, and the scores in both lists are added together to obtain a total score. The solution with the highest total score is selected as the output.
@@ IV-D1 Exact evaluation
When $S$ is fast recognized as infeasible solutions, it takes $O(1)$. Otherwise, in order to verify whether $S$ is feasible, the maximum number of iterations that the evaluation process is $T=L^{m}(1-P_{0})-l_{min}^{m}$, where $l_{min}$ is the index that the combination of $l_{min}$ samples of solution is larger than $T_{\max}$. The maximum size of heap is $k_{0}+(m-1)T=\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+(m-1)L^{m}(1-P_{0})-(m-1)l_{min}^{m}=O(mT)$. At each iteration, the evaluation approach needs to sort the heap once, so the total number of evaluation times is $O(\log(mT))=O(\log m+\log T)=O(\log T),m\ll L$. The total time complexity is $O(mL\log L+\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+T\log T)=O(\max\{T\log T,ml_{min}^{m-1},m^{2}l_{min}^{m-2},...,ml_{min}\})$.
@@ IV-D2 Accelerated Monte Carlo evaluation
The complexity of the accelerated Monte Carlo evaluation is $O((|G|+\alpha\times MonteCarlo)\times Eval)$, where $MonteCarlo$ denotes the simulation times of Monte Carlo simulations and $|G|$ is the size of the filtered list. If $S$ is infeasible recognized by filtered list, $\alpha=0$, otherwise $\alpha=1$.
@@ IV-D3 DDALS
  • CP includes sorting and solution evaluation. The times of evaluation are up to $m\times N$ Thus, the complexity is $O(Eval\times mN)$, where $N=\max|N_{i}|$.
  • LSS requires up to $m\times N$ times of evaluation to search for a better solution. The complexity is $O(Eval\times mN)$.
  • For Degrade procedure, the worst case is that only one combination of items is feasible. Thus, the complexity is $O(Eval\times mN)$.
  • FSS iterates through a maximum of combinations of classes $\binom{m}{2}\times N^{2}$. The complexity is $O(Eval\times m(m-1)N^{2}/2)=O(Eval\times m^{2}N^{2})$.
Note that SFE only takes $O(1)$. Therefore, the total time complexity of DDALS is $O(Eval\times mN+MaxIter\times(Eval\times mN+Eval\times mN)+Eval\times N^{2}m^{2}+1)=O(Eval\times\max\{N^{2}m^{2},MaxIter\times mN\})$.
@@ V Benchmark Design
As CCMCKP has not been studied before, there is a lack of corresponding benchmarks. To evaluate the performance of algorithms for solving CCMCKP, we designed two sets of benchmark with different problem scales and the intensity of chance constraints.
@@
The first set aims to simulate the scenarios where the delay of items follows different common continuous distributions, named LAB. The distributions of items are randomly chosen from uniform, truncated normal, fatigue-life, bimodal and gamma distributions. The mean of items is uniformly distributed in $U(2,8)$, where $U(2,8)$ denotes a uniform distribution over the interval [2,8]. In the case of gamma distribution, the mean value is $\mu\sim U(0.5,2.5)$ and the variance is $\sigma^{2}\sim U(0.05,0.625)$. The cost, representing the value associated with taking each item, is randomly generated from $U(1,10)$. This corresponding to an uncorrelated situation in the widely used benchmark of classic MCKP [22]. The probability density functions for different distributions in LAB are shown in Figure 4a.
@@
The second set, denoted as APP, is designed to simulate the real-world business requirements of the 5G E2E network from a certain telecommunications company. In APP, the distributions of item are also randomly chosen in the set of uniform, truncated normal, fatigue-life, bimodal and gamma distribution. However, in the 5G E2E network, each data package transmission is restricted to a time window with a fixed size, indicating that packets must be transmitted successfully within a set interval. This distinctive feature is known as retransmission mechanism [15]. Retransmission mechanism implies that each item in all classes has a probability of being retransmitted when a transmission failure occurs. In APP, the first attempt to transmit a data package must be completed within $(0,10]$ ms with a success probability of 0.9. If the initial attempt fails, the second attempt occurs with a delay of $(10,20]$ ms and a success probability of 0.9. This pattern continues for the third and fourth retransmissions. Finally, the transmission is recognized as a failure if the attempt is more than four times. The time delay for an item to transmit a data package falls within the intervals $(0,10]$, $(10,20]$, $(20,30]$, $(30,40]$ with probabilities of $0.9,0.09,0.009,0.001$, respectively. The distributions of items involved in APP are illustrated in Figure 4b. The cost generation formula is defined as $c_{ij}=10/(\mu_{ij}+\sigma_{ij})\times U(0.8,1.2)$. This formulation is designed to induce a decrease in cost-effectiveness as the time delay decreases, simulating the marginal diminishing effect of cost-effectiveness in practical scenarios.
For instances in LAB and APP, we set the values of $m$ and $N$ similar to [22]. Additionally, we categorize them into small-scale and large-scale instances. The sample size is set as 30 for small-scale instances and 500 for large-scale instances. The settings of all instances are shown in Table I. “11,14” of $T_{\max}^{LAB}$ means that $T_{\max}$ of LAB is set 11 or 14 as two different intensity of chance constraints.
TABLE I: Instances setting
Instance $m$ $N$ $L$ $T_{\max}^{LAB}$ $T_{\max}^{APP}$